OpenStudy (anonymous):

What is the equation of the line that passes through the point (4, 1) and is perpendicular to the line y = –2x + 4?

OpenStudy (mertsj):

Do you know point-slope form of the equation of a line?

OpenStudy (anonymous):

no i dont know any algebra im trying to help my sister finish her assignments. shes in the hospital. i dont need anyone to teach me anything i just dont want her to fail. any good answers anyone has is greatly appreciated

OpenStudy (mertsj):

Sorry. It is contrary to the code of conduct to just post answers. Too many kids try to avoid learning that way.

OpenStudy (anonymous):

could u show me how to work it out then? dont want anyone to get in trouble.

OpenStudy (mertsj):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

First, determine the slope of the given line. x + 2y = 6 Solve for y to get this in slope-intercept form. Remember to switch signs when you switch sides. 2y = -x + 6 Divide all terms on both sides by 2. y = -0.5 x + 3 So the slope of the given line is the coefficient of x, which is -0.5 = -1/2 The slope of a perpendicular line is the negative reciprocal of that slope, which would be 2. Use the point-slope form to build the equation with the point (4,2) y - y1 = m(x - x1), where (x1, y1) is a point on the line and m is the slope. y - 2 = 2(x - 4) This is an equation that satisfies the question. To get this into slope-intercept form, expand, simplify and solve for y. y = 2x - 8 + 2 y = 2x - 6

OpenStudy (mertsj):

That is point-slope form. The given line has slope -2 so the perpendicular has slope 1/2 replace y1with 1, m with 1/2 and x1 with 4 and simplify. That will be the equation of the desired line.

OpenStudy (anonymous):

Yep, take it with a little grain of salt

OpenStudy (anonymous):

so what do i have to simplify exactly?

OpenStudy (anonymous):